OpenStudy (deathlynine):

I will medal. Can someone show me the steps on how to answer this -4(x + 2) > 8

OpenStudy (stbest):

-4(x + 2) > 8 X has to be something that will make the full added number greater than 8 so lets try 5 5+2 =7 and 7 x -4 is -28 so its surly no that lets try a negative number -5 -5+2= -3 and -3x-4=12 and 12 is surely greater than 8 ()

OpenStudy (kashmoneyjr):

-4(x + 2) > 8 Distribute, so: -4x-8>8 Add 8 to both sides -4x>16 Now, you have to divide by -4. However, since it is negative, you have to switch from > to < Finally, we get x<-4 Basically, treat the inequality as if it had an = instead of < or >. The only time when you have to worry about < or > is when you multiply/divide by a negative number, which flips the sign.

OpenStudy (deathlynine):

Thanks both of you. If only Open Study allow me to medal more than once.